OpenStudy (anonymous):

v=At^3-Bt where t is time, identify the dimensions defined by A and B

OpenStudy (anonymous):

Basically need a way to define velocity by something time cubed - something time. trying to play around with initial velocity or initial acceleration - the other, but i'm not getting anything that makes any sense

OpenStudy (anonymous):

they can be scalar values. I dont understand what you want to ask

OpenStudy (anonymous):

I'm supposed to determine what dimensions A and B can be (like say it was N=m*A then A what be defining m/s^2) I just don't know how to define v using something - something else besides Vf-Vo

OpenStudy (anonymous):

V=m/s t=s so your one is \[V=A.t ^{3}-B.t\] \[m/s=A.s ^{3}-B.s\] the right part must be m/s so \[A.s ^{3}=m/s\] \[A=m/s ^{4}\] B.t=m/s \[B=m/s ^{2}\] that ?

OpenStudy (anonymous):

I'm going with that because that was the only way I could think to make it make sense, just reiterate the v by making A and B nonsensical just to force them to equal v. (make it non-applicable, but purely a definition)